Jane Miller has a house assessed at $45,000 (45 percent of its sale value). Her tax rate is $4.20 per $100 AV. Tom Gale has a house assessed at $55,000 (50 percent of its sale value). His tax rate is 35.25 mills. Who pays the greater tax? How much greater?. If both houses were taxed on full market (sale) value with a true tax rate of 23.3 mills, who would pay the greater tax?How much greater?
Answer #1

Answer :-

Tax Paid -

  • Jane Miller = Assessed value * Tax Rate

= $55,000 * 35.25 * 1/1000 (or 3.52)

= $1938.75

  • Tom Gale = Assessed value * Tax Rate = $45,000 * 4.2%   = 1,890
  • Jane Miller has to  pay more on property taxes in this case. Note:- The term millage rate is came from latin word ie., mill for short ,which means 'thousandth part (1/1000).Thus ,35.25 mills = 35.25*1/1000 ie.,..0.0352 or 3.52%